# The Essential Role of Causality in Foundation World Models for Embodied AI

### v1

@@ 4 Research opportunities
The foundation model for Embodied AI extends beyond traditional vision and language realms. Robotics necessitates the integration and interpretation of diverse sensory data, including tactile and various torque sensors, covering a broad range of modalities which vary from one agent to the next. This requirement surpasses the capabilities of current vision and language-based foundation models. Developing a system capable of effectively processing and utilizing these complex sensory data is crucial for enhancing robotic capabilities, especially in more intricate and nuanced tasks. On the other hand, building a causally-aware FWMs that comprehends high-dimensional observations from a large number of potential different modalities to model the underlying mechanisms is very challenging and presents multiple research opportunities:
@@
Secondly, agents vary widely in their sensor and effector configurations, from humanoid robots like Tesla’s Optimus to quadrupedal robots like Boston Dynamics’ Spot, to virtual cars, each with unique sensory inputs and available effectors. For FWMs to make precise action-outcome predictions, they must semantically interpret these diverse inputs and controls, necessitating a representational model that encapsulates the essence of an agent’s design to enable generalization across different configurations.
@@
As discussed, the canonical research in causality and RL shares the same objective but primarily differs in the main source of the data paradigm: observational data and intervention data. Traditionally, causality seeks to learn the world model using existing data, assuming that new interventions are not feasible, while RL learns through interactions. For embodied AI with FVWM, there will be access to a vast amount of offline data from humans, such as vision language data, human motion data, and robotics data, along with the possibility to obtain new intervention data. We need to utilize all different data paradigms in the most efficient way, breaking down the barriers between foundation models, RL, and causality. For example, we can look into debiasing the observational data using causal-aware methods such as CInA (Zhang et al., 2023c). Moreover, how to efficiently combine these algorithms and data paradigms in a single foundation model is also non-trivial with a lot of potentials. For example, if the model has a high loss of predicting how the dynamics change when a modality in the past is altered, then it should aim to answer this either through weighting the loss higher for the altered data point, or aim to collect this information through online interactions. This can be done in an iterative manner where a base model is trained from offline data, then allowed access to the online simulator to satisfy its curiosity (i.e. counterfactuals) and so on and so forth (Zeng et al., 2023; Gasse et al., 2021; Ivanova et al., 2023; Tigas et al., 2022; Seitzer et al., 2021; Sontakke et al., 2021). Such causally curious foundation model can improve the sample efficiency of training large models by reducing computing times, enabling multi-step reasoning and reduced hallucinations, and thereby no longer requiring RLHF (Casper et al., 2023) to finetune the trained models for specific tasks. A veridical causal foundation model can therefore be directly employed to make multi-step future predictions and directly enable planning methods (Zhang et al., 2023d) to find the sequence of the path leading to the desired outcomes. This is also consistent with the consideration of active data collection, where RL and experimental design are closely interlinked (González et al., 2016; Ma et al., 2019; Yin et al., 2020; Foster et al., 2021). Additionally, such an FVWM can massively accelerate planning methods to learn optimal policies in a sample efficient manner.
Empirical driven research and evaluation
In incorporating causality research into the foundation models for embodied AI, the key lies in the successful execution of actions, which depends on accurately predicting or imagining the consequences of different actions. In embodied AI research, physical meaningfulness is essential. Therefore, we believe that shifting the research culture towards being empirically driven, especially for causal researchers, and developing relatable, realistic, and economical evaluation methods is crucial for success.
@@
Evaluating real-world scenarios with a physical device is very costly in terms of time and potential risks, although it is necessary for the ultimate success of foundation embodied AI. Thus, constructing a curriculum of tasks that are easy to evaluate is key. This means evaluations can be performed at scale, replicated by any research institute, have clear metrics demonstrating progress, and provide a reliable approximation for end tasks in the real world. This requires not only further advancements in embodied AI in simulated environments like CausalWorld (Ahmed et al., 2020; Bauer et al., 2022) and Habitat 3.0 (Puig et al., 2023) but also the creation of innovative tasks that aid in this process. For example, evaluation on counterfactual prediction can be a good approximation for assessing true world understanding. However, we need to consider concrete tasks and evaluation metrics. In this case, the task could involve a generated video (Yin et al., 2023), while the metric needs to be more nuanced, as pixel space evaluation may have drawbacks, as discussed in computer vision research communities.
@@ 5 Deploying Embodied AI
Many factors affect how FVWMs can impact the deployment of Embodied AI in the real world. In this chapter, we focus on what we believe will be one of the most important factors, which is FVWMs ability to greatly scale up the deployments of Embodied AI in robotics, and thereby also robotics in general.
@@
General purpose robots form perhaps the most exciting prospect of generally capable Embodied AI. By compatibility with human environments and tools, robots of this kind could finally bridge the gap between the physical and digital worlds and allow large scale automation of physical tasks. Potential applications range from automating domestic tasks (Puig et al., 2023), nursing (Ohneberg et al., 2023), work in hazardous environments (Petereit et al., 2019) to manufacturing (Kheddar et al., 2019). We discuss ethical considerations regarding societal impact in our impact statement. Recent work has demonstrated zero-shot solutions to complex tasks, yet shows poor performance when faced with new environments. Research in causal FVWMs should prioritize consideration of diverse modalities and platforms, enabling success across a wide range of tasks and environments.
@@
While general purpose robots capable of zero-shot learning stand to reduce deployment time, specialized robots will remain more cost effective in some scenarios - especially for large-volume production. Examples include farming (Aravind et al., 2017), scientific lab exploration (Ren et al., 2023; Boiko et al., 2023) and industrial welding machines (Grau et al., 2020). Implementing specialized robots is a time-consuming process and with the learning to perform task without explicit programming is prioritized on the research agenda (Dzedzickis et al., 2022). However, current learning implementations are typically limited to a single environment (Meyes et al., 2017). We believe that causal-aware FVWMs may solve these constraints and can therefore dramatically impact the deployment of industrial specialized robots. An important testbed for this work will therefore be operation across robot morphologies specialized with different sensing and tooling attachments.
@@
Industrial robots have typically been confined to isolated cells. Through constraints such as force limitations, more collaborative modes of operation have been enabled, which have resulted in improvements in productivity (Dzedzickis et al., 2022). In other applications, such as nursing homes, direct physical interaction with humans is vital. Yet in such situations, the constraints from highly controlled industrial environments may not be sufficient to ensure physical safety. A thorough grasp of the outcomes resulting from physical interactions, facilitated by FVWMs, presents an opportunity to enhance deployment safety, provided that the evaluation steps outlined in Section 4 are meticulously followed.
@@ 6 Related Work
There have been a number of surveys on foundation models with a focus on recent research advances, challenges, and opportunities (Zhou et al., 2023a; Bommasani et al., 2021). Regarding embodied AI, Firoozi et al. (2023); Xiao et al. (2023) survey the role of pretrained foundation models in robotics for perception, navigation, decision making and control. Recent studies have explored the use of vision-language FMs in embodied AI tasks to improve sample efficiency and performance abilities (Khandelwal et al., 2022; Batra et al., 2020; Shridhar et al., 2022). Moreover, numerous surveys from the perspective of robotic tasks have been published, providing insights applicable to both deep learning and machine learning at large (Yin et al., 2021; Newbury et al., 2023). Recent research advances have increasingly focused on utilizing foundation models for embodied AI. A number of methods leverage large LLMs and VLMs for planning, re-planning and navigation (Lin et al., 2023a; Schumann et al., 2023; Skreta et al., 2024). Also, more recent work focuses on generalization across different environments and tasks for robots based on VLM, such as SayCan (Brohan et al., 2023b), RT-2 (Brohan et al., 2023a), GATO (Reed et al., 2022), AutoRT (Ahn et al., 2024), and autonomous driving, such as GAIA (Hu et al., 2023).
@@
In recent years, books (Pearl, 2009b; Hernán & Robins, 2010; Cunningham, 2021; Imbens & Rubin, 2015; Spirtes et al., 2000; Peters et al., 2017) as well as surveys (Spirtes & Zhang, 2016; Kitson et al., 2021; Shojaie & Fox, 2022; Zanga et al., 2022) have been published, drawing attention to causality and causal machine learning in general. Related to our work is the recent overview of open problems from (Kaddour et al., 2022), which connect different machine learning approaches to causality (Schölkopf, 2022); and a survey and position paper on causal reinforcement learning (Zeng et al., 2023) which aligns with our position of breaking the boundary of these field for advanced performance. There is large body of research work considering generic deep learning and causality including (Schölkopf et al., 2021; Jin et al., 2022; Zhou et al., 2023b; Berrevoets et al., 2023).
@@
World models, often used in RL, aim to model the transition dynamics which are expected to be intrinsically causal. It is a longstanding challenge, where early work has used standard sequence models for this purpose. For example, Ha & Schmidhuber (2018) employ a Variational Autoencoder (VAE) to condense visual inputs into a latent representation with a RNN to model the dynamics. Other studies have focused on learning latent state dynamics models, as seen in Hafner et al. (2019); Gelada et al. (2019), while advancements in the Dreamer series (Hafner et al., 2020, 2023), TransDreamer (Chen et al., 2022) and Micheli et al. (2022) highlight the benefits of discrete latent variables and transformers for latent dynamics. Recent works have also leveraged diffusion models directly for world models (Rigter et al., 2023; Anonymous, 2024a).
The community has increasingly recognized the important role of causal properties in world models. Recently, Anonymous (2024b) proved in theory that any agents capable of generalizing under domain shift must learn a causal model, and there have been many works trying to incorporate causality. Zhang et al. (2020b, 2019) use causal core sets (Zhang et al., 2020a) to model latent transition dynamics. Li et al. (2020); Lu et al. (2018), on the other hand, considers the confounding effect during the modelling of state transitions. Another line of work (Annabi, 2022; Volodin et al., 2020; Zhu et al., 2022; Wang et al., 2022; Tomar et al., 2021) explicitly build a SEM among variables in the states, allowing the transition dynamics to only depend on causally-relevant feature. Unfortunately, existing causal world models fall short as FWMs due to their inability to generalize across domains and tasks, and to process multi-modal inputs efficiently. Future research on causal-aware FVWM will bridge the gap and enable embodied AI.
@@ 7 Conclusion
This position paper, focusing on world models for Embodied AI, does not propose new technological advancements with direct societal consequences in isolation. However, it is crucial to address the broader social implications of the field. This is a topic of intense debate with a multitude of works predicting its outcome and potential mitigation. However, a reoccurring theme in the debate is that the whole of society to ensure an equitable transition to rapidly increasing automation. Willcocks (2020) argues that such a transition will unlikely result in a net job loss, and will be vital to mitigate changing demographics which are predicted to lead to labour shortages. Using foundation models in embodied AI shifts the focus of automation from tasks that are dirty, dull, and dangerous (3D) to those that necessitate advanced cognitive and reasoning skills. As a result, the implications of widespread deployment in automating various job types could vary significantly, necessitating careful consideration by society.
Another important consideration is that Embodied AI should be ethical and fair. For the real-world evaluation of embodied AI, safety and ethical considerations are key, as there is the possibility of inducing physical harm. Moreover, Causal FVWMs may reduce biases and lead to fairer algorithms (Kusner et al., 2017; Zhang & Bareinboim, 2018). However, as suggested by Nilforoshan et al. (2022) may not directly translate to equitable decision making. Instead it will remain important for researchers to develop this area further.
@@ Acknowledgment
The authors wish to express gratitude to Katja Hofmann, Nan Duan, Anastasia Varava, Xin Tong, Tom Minka, Jiang Bian, David Sweeney and Baining Guo, especially to John Langford, whose discussions have inspired and aided the manuscript.

### v2

@@ 4 Research opportunities
The foundation model for Embodied AI extends beyond traditional vision and language realms. Robotics necessitates the integration and interpretation of diverse sensory data, including tactile and torque sensors, covering a broad range of modalities which vary from one agent to the next. This requirement surpasses the capabilities of current vision and language-based foundation models. Developing a system capable of effectively processing and utilizing these complex sensory data is crucial for enhancing robotic capabilities, especially in more intricate and nuanced tasks. On the other hand, building a causally-aware FWMs that comprehends high-dimensional observations from a large number of potential different modalities to model the underlying mechanisms is very challenging and presents multiple research opportunities.
@@
Second, agents vary widely in their sensor and effector configurations, from humanoid robots like Tesla’s Optimus to quadrupedal robots like Boston Dynamics’ Spot, to virtual cars, each with unique sensory inputs and available effectors. For FWMs to make precise action-outcome predictions, they must semantically interpret these diverse inputs and controls, necessitating a representational model that encapsulates the essence of an agent’s design to enable generalization across different configurations.
@@
As discussed, canonical research in causality and RL shares similar objectives but differs in the source of the data: observational data versus interventional. Traditional causality research seeks to learn the world model using existing data, assuming that new interventions are not feasible, while online RL learns through interactions. For Embodied AI with FVWM, there will be access to a vast amount of offline observational data from humans, such as vision language data, human motion data, and robotics data, along with the possibility to obtain new intervention data. We need to utilize all different data paradigms, breaking down the barriers between foundation models, RL, and causality.
Pre-training FVWMs from offline data alone may be sufficient to obtain good performance in some scenarios. However, we anticipate that gathering additional online interaction data (Lee et al., 2022) will be necessary to obtain robust performance across general tasks and environments. This data should be collected in a way such that it enables the model to reason about interventional and counterfactual scenarios and therefore integrating different levels of causal reasoning (Pearl, 2009b). Therefore, a key research direction will be developing algorithms that efficiently gather the most useful intervention data online, enabling fast adaptation:
  • The related areas of active learning and experimental design selectively query datapoints to maximize information gain (González et al., 2016; Ma et al., 2019; Gong et al., 2019; Yin et al., 2020). Similar ideas have been explored in the context of training world models by collecting data where the world model has the greatest epistemic uncertainty (Sekar et al., 2020; Rigter et al., 2024). However, maximising uncertainty does not necessarily lead to the collection of data that is most useful for improving decision-making (Pathak et al., 2017). Another paradigm then is to incentivize agents to collect semantically diverse data from online interactions based on representations grounded in natural language (Gupta et al., 2022; Tam et al., 2022) or control-relevant information contained in observations (Lamb et al., 2022; Islam et al., 2022).
  • The research into unsupervised environment design (Dennis et al., 2020; Samvelyan et al., 2023) aims to learn to adapt the difficulty of the environment to the agent’s learning progress and can produce a natural curriculum of increasingly complex environments for agents to interact with and collect data from, thereby improving the knowledge of world model in a continual manner.
In summary, by training a causally aware model, FVWMs may unlock opportunities to develop algorithms that efficiently gather the intervention data that is most relevant for improving decision-making. Such an approach has the potential to greatly expand the generality and reliability of the world model while reducing the cost of online data acquisition.
Improving planning
FVWMs aim to facilitate long-term decision-making in Embodied AI systems. It is well-known that improved prediction accuracy leads to improved planning (Kearns & Singh, 2002; Janner et al., 2019). However, even with an accurate world model, optimising behaviour with large prediction models is still a challenging problem and requires either reinforcement learning (Hafner et al., 2020), search (Schrittwieser et al., 2020), or sophisticated model-predictive control (Hansen et al., 2022) within the world model. While these approaches have proven successful in narrow tasks, it is unclear whether they will scale to foundation world models trained across diverse tasks and environments. By modelling causal structure, FVWMs have the potential to learn representations that are highly suitable for planning and optimisation. Thus, investigating how the representations learnt by FVWMs can improve not only prediction accuracy and generalization, but also planning, is another key research direction.
Latent Dynamic Representations for World Models
To enable sample efficient learning and planning with strong generalization abilities and causal reasoning, it will be crucial for FVWMs to learn suitable abstract representations of the world. Current large world models such as GAIA (Hu et al., 2023), LWM (Liu et al., 2024) learn to encode the entire images into a latent space based on a reconstruction loss which can capture many irrelevant details for decision-making. Therefore, minimal state representations (Lamb et al., 2022; Islam et al., 2022), which only capture agent controllable features of the environment, are more suitable for causal reasoning and planning. Some possible approaches are using regularisation techniques such as sparsity or causal priors (Li & Han, 2024; Killedar et al., 2021) in the latent space, which can be thought of as encouraging the latent space to learn the SEM (Rajendran et al., 2024). Other possibilities include changing the model architecture to learn latent representations that disentangle task-irrelevant visual details from task-relevant details and dynamically adapt these representations as task distribution of interest changes. With improved state representations, we expect that the model should be able to capture causal relationships and therefore learn in a more sample-efficient way and generalise better.
Empirically-driven research and evaluation
The goal of Embodied AI research is to improve the performance and applicability of embodied agents in real-world scenarios. Therefore, to measure success we should focus on empirical performance on metrics related to physically meaningful tasks, rather than theoretical results. Thus, shifting the research culture towards being empirically-driven, especially for causal researchers, and developing relatable, realistic, and economical evaluation methods is crucial for success.
@@
Evaluating real-world scenarios with a physical system is very costly in terms of time and potential risks, although it is necessary for the ultimate success of Embodied AI. Thus, constructing a curriculum of tasks that are easy to evaluate is key. This means evaluations can be performed at scale, replicated by any research institute, have clear metrics demonstrating progress, and provide a reliable approximation for end tasks in the real world. This requires not only further advancements in Embodied AI in simulated environments like CausalWorld (Ahmed et al., 2020; Bauer et al., 2022) and Habitat 3.0 (Puig et al., 2023) but also the creation of innovative tasks that aid in this process. For example, evaluation on counterfactual prediction can be a good approximation for assessing true world understanding. However, we need to consider concrete tasks and evaluation metrics. In the case of counterfactual prediction, the task could involve generating a video (Yin et al., 2023), while the metric needs to be nuanced as pixel-space evaluation has drawbacks (Fernandez-Moral et al., 2018).
@@ 5 Deploying Embodied AI
Many factors affect how FVWMs can impact the deployment of Embodied AI in the real world. In this chapter, we focus on what we believe will be one of the most important factors, which is FVWMs ability to greatly scale up the deployments of Embodied AI in robotics.
@@
General purpose robots form perhaps the most exciting prospect of generally capable Embodied AI. By compatibility with human environments and tools, robots of this kind could finally bridge the gap between the physical and digital worlds and allow large-scale automation of physical tasks. Potential applications include automating domestic tasks (Puig et al., 2023), nursing (Ohneberg et al., 2023), work in hazardous environments (Petereit et al., 2019), and manufacturing (Kheddar et al., 2019). We discuss ethical considerations regarding societal impact in our impact statement. Recent work has demonstrated zero-shot solutions to complex tasks, yet shows poor performance when faced with new environments. Research in causal FVWMs should prioritize consideration of diverse modalities and platforms, enabling success across a wide range of tasks and environments.
@@
While general purpose robots capable of zero-shot learning stand to reduce deployment time, specialized robots will remain more cost effective in some scenarios - especially for large-volume production. Examples include farming (Aravind et al., 2017), scientific lab exploration (Ren et al., 2023; Boiko et al., 2023) and industrial welding machines (Grau et al., 2020). Implementing specialized robots is a time-consuming process and therefore learning to perform tasks without explicit programming an important research area (Dzedzickis et al., 2022). However, current learning implementations are typically limited to a single environment (Meyes et al., 2017). We believe that causal-aware FVWMs may resolve these constraints and therefore reduce the cost to deploy specialized industrial robots. An important testbed for this work will therefore be operation across robot morphologies with different sensing and tooling attachments.
@@
Industrial robots have typically been confined to isolated cells. However, constraints such as force limitations have enabled more collaborative operation, improving productivity (Dzedzickis et al., 2022). In other applications, such as nursing homes, direct physical interaction with humans is vital. Yet in such situations, simple threshold constraints may not be sufficient to ensure physical safety while still allowing the robot to perform the tasks of interest. A thorough grasp of the outcomes resulting from physical interactions, as facilitated by FVWMs, presents an opportunity to enhance safety by developing more nuanced safety constraints.
@@ 6 Related Work
There have been a number of surveys on foundation models with a focus on recent research advances, challenges, and opportunities (Zhou et al., 2023a; Bommasani et al., 2021). Regarding Embodied AI, Firoozi et al. (2023); Xiao et al. (2023) survey the role of pretrained foundation models in robotics for perception, navigation, decision making and control. Recent studies have explored the use of vision-language FMs in Embodied AI tasks to improve sample efficiency and performance abilities (Khandelwal et al., 2022; Batra et al., 2020; Shridhar et al., 2022). Moreover, numerous surveys from the perspective of robotic tasks have been published, providing insights applicable to both deep learning and machine learning at large (Yin et al., 2021; Newbury et al., 2023). Recent research advances have increasingly focused on utilizing foundation models for Embodied AI. A number of methods leverage large LLMs and VLMs for planning, re-planning and navigation (Lin et al., 2023a; Schumann et al., 2023; Skreta et al., 2024). Also, more recent work focuses on generalization across different environments and tasks for robots based on VLMs, such as SayCan (Brohan et al., 2023b), RT-2 (Brohan et al., 2023a), GATO (Reed et al., 2022), AutoRT (Ahn et al., 2024), and autonomous driving, such as GAIA (Hu et al., 2023).
@@
In recent years, books (Pearl, 2009b; Hernán & Robins, 2010; Cunningham, 2021; Imbens & Rubin, 2015; Spirtes et al., 2000; Peters et al., 2017) as well as surveys (Spirtes & Zhang, 2016; Kitson et al., 2021; Shojaie & Fox, 2022; Zanga et al., 2022) have been published, drawing attention to causality and causal machine learning in general. Related to our work is the recent overview of open problems from (Kaddour et al., 2022), which connect different machine learning approaches to causality (Schölkopf, 2022); and a survey and position paper on causal reinforcement learning (Zeng et al., 2023) which aligns with our position of breaking the boundary of between these fields. There is large body of research work considering generic deep learning and causality including (Schölkopf et al., 2021; Jin et al., 2022; Zhou et al., 2023b; Berrevoets et al., 2023).
@@
World models, often used in RL, aim to model the transition dynamics which are expected to be intrinsically causal. It is a longstanding challenge, where early work has used standard sequence models for this purpose (Schmidhuber, 1990). More recent research has focused on using Variational Autoencoders (VAEs) to map inputs into a latent representation to enable dynamics modelling with high-dimensional observations (Ha & Schmidhuber, 2018; Hafner et al., 2019; Gelada et al., 2019; Schrittwieser et al., 2020). Advancements in the Dreamer series (Hafner et al., 2020, 2023) and transformer-based world models (Chen et al., 2022; Micheli et al., 2023; Robine et al., 2023) highlight the benefits of discrete latent variables and transformers for latent dynamics. Recent works have also leveraged diffusion models for world models (Rigter et al., 2023; Alonso et al., 2023) to improve prediction accuracy at the cost of increased computation.
The community is increasingly recognizing the important role of causal properties in world models. Recently, Richens & Everitt (2024) proved that any agent capable of generalizing under domain shifts must learn a causal model, and there have been many works trying to explicitly incorporate causality. Zhang et al. (2020b, 2019) use causal core sets (Zhang et al., 2020a) to model latent transition dynamics. Li et al. (2020); Lu et al. (2018), on the other hand, considers the confounding effect during the modelling of state transitions. Another line of work (Annabi, 2022; Volodin et al., 2020; Zhu et al., 2022; Wang et al., 2022; Tomar et al., 2021) explicitly build a SEM among variables in the states, allowing the transition dynamics to only depend on causally-relevant feature. Unfortunately, existing causal world models fall short as FWMs due to their inability to generalize across domains and tasks, and to process multi-modal inputs efficiently. Future research on causal-aware FVWM will bridge the gap and enable Embodied AI.
@@ 7 Conclusion
This position paper, focusing on world models for Embodied AI, does not propose new technological advancements with direct societal consequences in isolation. However, it is crucial to address the broader social implications of the field. This is a topic of intense debate with a multitude of works predicting its outcome and proposing how to mitigate risks. However, a reoccurring theme in the debate is that is will be necessary to ensure an equitable transition of society to rapidly increasing automation. Willcocks (2020) argues that such a transition will unlikely result in a net job loss, and will be vital to mitigate changing demographics which are predicted to lead to labour shortages. Using foundation models in Embodied AI shifts the focus of automation from tasks that are dirty, dull, and dangerous (3D) to those that necessitate advanced cognitive and reasoning skills. As a result, the implications of widespread deployment in automating various job types could vary significantly, necessitating careful consideration by society.
Another important consideration is that Embodied AI should be ethical and fair. For the real-world evaluation of embodied AI, safety and ethical considerations are key, as there is the possibility of inducing physical harm. Moreover, Causal FVWMs may reduce biases and lead to fairer algorithms (Kusner et al., 2017; Zhang & Bareinboim, 2018). However, as suggested by Nilforoshan et al. (2022), this may not directly translate to equitable decision making. Instead it will remain important for researchers to develop this area further.
@@ Acknowledgment
The authors wish to express gratitude to Katja Hofmann, Jan Peters, Nan Duan, Anastasia Varava, Xin Tong, Tom Minka, Jiang Bian, David Sweeney Baining Guo and John Langford, whose discussions have inspired and aided the manuscript.